### v2

# Data-driven Discovery for Robust Optimization of Semiconductor Nanowire Lasers

Abstract
Active wavelength-scale optoelectronic components are widely used in photonic integrated circuitry, however coherent sources of light – namely optical lasers – remain the most challenging component to integrate. Semiconductor nanowire lasers represent a flexible class of light source where each nanowire is both gain material and cavity; however, strong coupling between these properties and the performance leads to inhomogeneity across the population. While this has been studied and optimized for individual material systems, no architecture-wide insight is available. Here, nine nanowire laser material systems are studied and compared using 55,516 nanowire lasers to provide statistically robust insight into performance. These results demonstrate that, while it may be important to optimise internal quantum efficiency for certain materials, cavity effects are always critical. Our study provides a roadmap to optimize the performance of nanowire lasers made from any material: this can be achieved by ensuring a narrow spread of lengths and end-facet reflectivities.
keywords:
high-throughput, nanowire lasers, interferometry, photoluminescenceUoM] Department of Physics and Astronomy and The Photon Science Institute, The University of Manchester, M13 9PL, United Kingdom Jena] Institute for Solid State Physics, Friedrich Schiller University Jena, Max-Wien-Platz 1, 07743 Jena, Germany ANU] Department of Electronic Materials Engineering, Research School of Physics, The Australian National University, Canberra, ACT, 2601 Australia Zhejiang] College of Integrated Circuits, Zhejiang University, Hangzhou, Zhejiang 311200, China Cambridge] Department of Engineering, University of Cambridge, CB2 1PZ, United Kingdom ANU] Department of Electronic Materials Engineering, Research School of Physics, The Australian National University, Canberra, ACT, 2601 Australia UCL] Department of Electronic and Electrical Engineering, University College London, WC1E 7JE, United Kingdom Cambridge] Department of Engineering, University of Cambridge, CB2 1PZ, United Kingdom Jena] Institute for Solid State Physics, Friedrich Schiller University Jena, Max-Wien-Platz 1, 07743 Jena, Germany UoM] Department of Physics and Astronomy and The Photon Science Institute, The University of Manchester, M13 9PL, United Kingdom
The miniaturisation of opto-electronic devices is an area of rapid development across multiple research fields 1, 2, 3. In its simplest form this involves the development of devices at the centi- and millimeter-scale, enabling previously bulky apparatus to become light-weight and portable, facilitating accurate field-based testing, with applications including quality control 4 and medical screening 5. Further miniaturisation, to the micron-scale, has yielded significant advancements in commercial technology. This includes micro-LEDs 6, which form the foundation of ultra-efficient, high-resolution and wide colour-gamut display technologies 7, enabling commercial virtual and augmented reality headsets 3.
Nanoscale coherent light sources are widely sought after as biological probes8, 9, sensors, in quantum applications 10, 11 and most often as active components for photonic integrated circuitry12, 13, 2, 14 that integrate micron-scale light sources, waveguides and detectors at high densities to perform light-based computing processes 15. A wide range of architectures have been proposed as solutions, spanning off-chip coupling, flip-chip integration and hetero-epitaxial growth16, 17, 18. However, repeatable, high-density and high-yield laser sources are rare. Of particular concern is the strong coupling between material quality, cavity design and performance, which mean that small variations in defect density, surface passivation, hetero-junction parameters and geometry lead to significant effects for performance6, 20, 21.
Nanoscale lasers based on semiconductor nanowires (NWs) have long been proposed as coherent light sources that are well suited to integration22, 23, 24, 25, 26; these provide diverse emission properties, such as wavelength, pulse duration and modal structure, arising from the wide choice of material system, while being united in their fundamental operating principle of the nanowire forming a monolithic cavity and gain material. These NW lasers (NWLs) typically have diameters of 100’s nm and lengths up to 10’s m. High refractive indices result in strong optical confinement, and the narrow NW diameters often lead to a deterministic relationship between cavity, material and performance properties that necessitate optimization of each parameter 27. The variety of material systems has led to a divergence in this optimization and development, and to-date high quality systems have been demonstrated in chalcogenides28, 29, nitrides30, 31, perovskites32, 33, metal oxides34, 35 and III-Vs15, 37, 4. This range of systems provides significant advantages for selecting lasing wavelength, for high-speed operation39 or providing specific non-linear properties40. However, it is difficult to robustly compare the performance of each material system due to differences in excitation conditions and characterization methodologies (for example: laser wavelength, pulse duration and excitation spot size 41, 4, 5). This is exacerbated by the differences between growth mechanisms (for instance, selective-area growth43 versus vapour-liquid-solid44) and the inclusion of plasmonic39 or dielectric-enhanced34 emission.
Tight control of the device fabrication is therefore required to minimise variation in performance and to maximise the useful yield. This is a crucial step towards commercialisation of these miniature technologies, with yield requirements as high as, for example, in the microLED industry 45. Whilst an increasing number of reports now provide statistically robust intra-type results46, 15, 27, 47, comparison between nanowire types is rare, causing two significant roadblocks to further nanowire laser development: the lack of a common measurement methodology27 to produce comparable data, and the resultant impossibility of obtaining globally meaningful insight from specific studies.
In this work, we use the principles of experimental data-led discovery 48 to generate an experimental understanding of nanowire laser (NWL) performance, by studying nine different NWL designs with six independent automated experiments, which are detailed in ref15. This includes bright field optical imaging, low power continuous-wave photoluminescence (PL), fluence-dependent PL, emission imaging above the lasing threshold, time-correlated single photon counting (TCSPC) and interferometry, with 55,516 NWLs investigated in total. The approach enables the best-in-class NWLs to be identified, a statistically rigorous comparison of the performance of each NWL type and the identification of global trends and important factors that are independent of the NWL design.
A summary of the measurements performed on one individual InP NW are displayed in Figure 1. An example low power PL measurement is given in the supporting information. For each measurement, the NW was excited using the output from a PHAROS-ORPHEUS ultrafast laser-amplifier system with a repetition rate of and a wavelength of . The laser excitation spot was defocussed to uniformly excite the entire NW to achieve maximum gain. Further details of the experimental arrangement can be found in the supporting information and in ref 15. Power-dependent photoluminescence measurements exhibit a broad spontaneous emission peak at all pumping fluences, along with a narrow-band emission peak, above a threshold fluence, that is attributed to stimulated emission. The threshold is identified by analysing the light-in-light-out (LILO) curves of the excitation fluence vs the intensity of the emission, an example of which is shown in Figure 1(b). These curves are produced and processed using an established algorithm 15. In short, the emission intensity increases approximately linearly above lasing threshold, and a straight line was fit to this data and extrapolated to in order to estimate a lasing threshold fluence. Using this approach, the threshold for this NW was determined to be .
Optical imaging measurements for this NW are shown inset in Figure 1(b). Bright field images were used to measure the length of the NW, in this case. The diameter of the NW was also estimated using these images and found to be . It is important to note however that this width is similar to the optical resolution of the microscope used ( ), and is therefore likely an over-estimation. This limitation is discussed further in the supporting information.
TCSPC decays measured below the lasing threshold for the example NW are shown in Figure 1(c). For this NW, the 1/e decay time was at low fluence, increasing to at a higher fluence. These measured decay times are likely representative of a combination of non-radiative and radiative recombination processes, and as the lifetime increases with increasing fluence, we are unlikely to be in an excitation regime where Auger-Meitner recombination plays an important role for this NW. We therefore suggest that the non-radiative process is related to point defects in the InP and that this fast recombination pathway becomes saturated at higher fluences, thus increasing the effective decay time. This is in agreement with earlier studies, where InP NWs, and SAG grown NWs, have shown low surface recombination (SR) velocities 49. However, it is important to note that the fluence-dependence varies on a NW-to-NW basis, and thus Auger-Meitner recombination will play some role at high fluences for a subset of the NWs studied. This is discussed further in the supporting information.
Finally, a time-gated interferogram of the lasing emission above threshold was measured, shown in Figure 1(d), using an approach detailed in ref 15. A modulation envelope can clearly be observed in the data. The laser coherence length of the radiation was defined as the full width at half maximum (FWHM) of this envelope, which was found to be for this NW. This coherence length is several orders of magnitude longer than the NW length (). This is typical behaviour for the strong waveguides that are formed by NWLs 50, and is comparable to previous measurements on alternative materials 15. Using this result, the end-facet reflectivity was calculated to be using a previously detailed approach 12. This reflectivity is enhanced beyond the planar Fresnel reflection coefficient for this material () 52, which may initially be surprising as laser modes extending outside of the small NW cross-section would be anticipated to reduce the reflectivity. However, previous work 50, 53 has shown that strong waveguiding effects for certain transverse modes (TE01 and TM01) can result in a reflectivity enhancement comparable to this measurement, for nanowire diameters greater than . This is compatible with the diameters of the NWLs studied in this work.
A large, multidimensional dataset was generated by performing the measurements in Figure 1 for NWs of different architectures, which are summarized in Table 1. This includes “bulk” selective-area-grown InP 3, vapour-liquid-solid grown surface-passivated GaAs 4, Zn-doped GaAs 1, core-shell quantum well (QW) heterostructures AlGaAs/GaAs 5 and GaAsP/GaAs, which has 2 growth batches - 1 and - 2 11, 12. These growth batches nominally have the same core/shell structure, and differ only in the duration of the axial NW growth, with batch 1 containing longer NWs on average. Also included are NWs which demonstrate excitonic emission at room temperature due to large exciton binding energies: CdS 8, GaN 13, and ZnO 10, 14. The NWs were investigated using the same experimental setup as the InP NW in Figure 1, and the excitation wavelength was changed depending on the sample. For the “bulk” and QW samples was used, for the CdS NWs, was used and for GaN and ZnO, was used. A full breakdown of the number of measurements on each NW type is included in the supporting information.
Table 1 shows an exemplary bright field optical image for each type of NWL. For comparison, images of the spatial emission profile above lasing threshold for the same wires are also shown in Table 1. In this regime, the majority of the light couples out of the end facets of the NW Fabry-Perot cavity. Each end facet acts as an independent, coherent, point source of light and interference fringes between these two sources are observed in the far field 4.
Using this dataset, it is trivial to identify the best-in-type lasers by assessing the lasing thresholds. The LILO curves for these champion NWLs are shown in Table 1. The lowest observed threshold was from a ZnO NWL. This is consistent with measurements using similar excitation conditions on NWs of comparable size and geometry 61. However, this is two orders of magnitude higher than the lowest-reported thresholds in the literature 62, highlighting the sensitivity of this measurement to the experimental conditions and NW size.
The GaAsP/GaAs - 1 NWLs also perform favourably, with a lowest threshold of , which is similar to the best in the literature for this material 11, 12. The alternative growth batch GaAsP/GaAs - 2, has a champion threshold which is more than five times higher - emphasising the impact that NW length can have on performance. Despite this, it is interesting to note that the best-in-class thresholds are very similar for all types of ”bulk” NWLs studied. This implies that, for the best-performing ”bulk” NWLs, the functional performance in the pulsed excitation regime may be limited by the same effects, independent of the material used. This is possibly because the excitation and lasing occurs on timescales faster than many material-dependent carrier recombination processes 12.
The lasing emission spectrum of each best-in-type NW is also shown in Table 1, where a variety of behaviours are observed. Emission from multiple longitudinal lasing modes are observed from the InP,
GaAsP/GaAs - 1 and CdS example NWs - where the mode spacing is related to the length of the NW cavity [ref]. In contrast, single mode lasing (SML) is observed for the exemplary GaAs, Zn:GaAs, AlGaAs/GaAs and GaAsP/GaAs - 2 NWs at fluences approximately twice that of the threshold fluence. SML is often desired due to the resultant high-spectral purity and is often the result of strong mode selectivity in NWs 63. his agrees with Table 1, where SML is mainly observed in the shorter NW examples. It should however be noted that SML examples can be found for each NW type.
The data-led approach enables our study to extend from simple best-in-type measurements (Figure 1), through providing robust intra-type measurements of single NWL types (Table 1), to understanding global (inter-type) NWL architecture effects. The lasing threshold fluence and the lasing yield (the percentage of NWs with a lasing threshold below their damage threshold) are displayed in Figure 2(a). The lowest median threshold is seen in ZnO, with . Furthermore, the GaAsP/GaAs - 1 NWLs, despite having one of the lowest threshold champion devices, demonstrate a large variation in the threshold, such that the median is , and close to that of the other NWLs. In contrast, the lasing yield for GaAsP/GaAs - 1 NWLs is one of the highest available (), and is significantly higher than the yield of ZnO NWLs (). Clearly the interplay between yield and threshold is not straightforward, and whilst best-in-class devices can offer insights into outlier performance, a statistical approach is required for repeatable inter-class comparison.
The largest median thresholds and threshold variation are seen in the AlGaAs/GaAs and GaAsP/GaAs - 2 core/shell QW NWLs, which have thresholds between 1000 and . These NWs also have the lowest yields, just 2 and ; this is likely because the lasing thresholds are close to the damage threshold of the NWs. Careful NWL heterostructure design is essential to ensure that the QWs overlap well with the transverse lasing modes, and this must be maintained across a population through growth optimisation.
The distribution of the lasing wavelength for each class is shown in Figure 2(b). These wavelengths depend upon both the laser gain spectrum and the cavity 4. The wavelength variation in each class can be as high as : as a result, the full NW population has continuous wavelength coverage across the spectral window between 700 and . While this variability presents the most significant obstacle to scale-up, this does provides a means of effectively tuning the wavelength by selecting a NWL with the desired properties64. To demonstrate this coverage, the lasing spectra for 12 NWLs are shown in Figure 2(b); these NWLs have been chosen to be single-mode and to change the lasing wavelength by , although it is possible to select NWLs based on wavelength with nm-level precision, due to the size of the NW population. This type of data-driven screening is useful as a way to target ideal devices for transfer printing into photonic integrated circuits 47.
The distribution of NW lengths are shown in Figure 2(c). The lengths vary between , which is the resolution of the optical imaging, and . The measured NW diameters were also found to largely correlate with the lengths: the distributions of which are shown in the supporting information. These distributions offer a simple assessment of the uniformity of the growth dynamics and the NWL transfer process 65. In this case, the InP NWs stand out as being the most uniform, with a length interquartile range (IQR) of ; as the only selective-area grown nanowires, this uniformity is a demonstration of the advantage of catalyst-free growth for geometrical homogeneity. The GaAsP/GaAs - 1 length distribution is also distinctive, as this has the largest median length of , which is more than longer than the other heterostructure NW types. This is a strong indication of the importance of the NW length on the performance: longer NWs have a larger volume of gain material and so higher overall gain when excited uniformly 6.
Figure 2(d) shows the distributions for the below-threshold photoluminescence lifetime measured from TCSPC decays. The InP, GaAs, AlGaAs/GaAs, GaAsP/GaAs - 2, CdS, GaN and ZnO NWs all exhibit lifetimes that are short when compared with radiative lifetimes of comparable materials 66, 67, 68, 69, 70, 71, with median values between 100 and , suggesting that non-radiative mechanisms play a role in carrier recombination.
SR may be anticipated to contribute to this non-radiative recombination, due to the large surface-to-volume ratio of the NWs, however, in many cases the impact of the surface is reduced. For example, InP NWs, grown to the same recipe, have been shown to have SR velocities that are orders of magnitudes below other III/V semiconductors 49. Furthermore, the GaAs NWs have a shell of AlGaAs which prevents carriers from reaching the surface 4, and a similar effect is achieved for the core/shell AlGaAs/GaAs 5 and GaAsP/GaAs NWs 11, 12, as the active QW region is isolated from the surface by the QW barriers. In contrast, the CdS, GaN and ZnO are not surface-passivated. The measurements in this paper do not directly address SR and therefore this cannot be ruled out as a pathway for non-radiative recombination in any of the NW samples.
The Zn-doped GaAs NWs have the fastest median lifetime, , which is limited by the minimum time resolution of the experiment. Previous studies on the same NWs observed rapid radiative recombination, because of the doping, with a lifetime of 3- 72, 1. This is sufficient to out-compete the SR in the NWs.
Interestingly, the two batches of GaAsP/GaAs NWs have contrasting lifetimes. Batch 1 has the longest median recombination lifetime of . This batch has been previously studied 15, where it was established that the recombination efficiency is around , meaning that the rates of non-radiative and radiative recombination are comparable. The median radiative lifetime is therefore , and the median non-radiative lifetime is . As the impact of SR is reduced in this sample, the non-radiative recombination is likely related to the material quality and the non-radiative defect density. The growth conditions for batch 2 were nominally the same as batch 1 and so it is surprising that the median lifetime is reduced to . This result suggests that the non-radiative rate, and hence the defect density, is higher in the second batch.
The end-facet reflectivity was calculated from measurements of the coherence length12 as shown in Figure 1(d), and the distributions are shown in Figure 2(e). This measurement was not possible on the NWs with emission wavelengths less than due to limitations in the equipment sensitivity at these wavelengths. The vast majority of NWs have reflectivities that are enhanced above the Fresnel reflection factors for planar surfaces. This is consistent with previous measurements of a single NW class 12, 15 and experimentally verifies the theoretical predictions that have underpinned the use of the NW architecture for two decades 50. The InP, GaAs, GaAsP/GaAs - 1 and CdS NWs show lower median reflectivities around 0.6 and IQRs up to 0.31. This large degree of variation may be associated with differences in NW diameter or transverse lasing mode - and thus changes in the waveguiding properties 50: irregular morphology at the end facets, either due to the NW transfer procedure 65, or from an abundance of lattice defects in this region 73, may also play a role. The other NWs, Zn:GaAs, AlGaAs/GaAs and GaAsP/GaAs - 2, have high median reflectivities between 0.89 and 0.91 and IQRs up to 0.08. While statistically significant, these findings are likely a result of selection bias in the measurements due to the low yield of these NWLs. Due to the nature of the measurement, reflectivities can only be determined for lasing NWs: thus, high material quality NWLs can operate with poor cavities, while low material quality (poor IQE) NWLs require a high quality cavity to show lasing below their damage threshold.
Whilst analysing each parameter distribution in isolation can provide useful information regarding inhomogeneity in each type of NW, a key benefit of the holistic high-throughput approach is the ability to directly compare the distributions of different parameters for the same NWLs and isolate correlations to investigate physical phenomena. In particular, the parameters that are most important in controlling the lasing threshold can be determined. This is facilitated by the large population sizes in this study: for example, if a correlative analysis is attempted in a small data study, with approximately 100 NWs, an important correlation between two parameters will be obscured by the significant variation in the other parameters, as suggested by the boxplots in Figure 2.
To investigate the most important factors that influence the lasing threshold, the datasets were normalised by subtracting the median and dividing by the standard deviation of each distribution for each NW type. The data was then grouped into 3 classes, as defined by Table 1: excitonic, “bulk” and heterostructure. This was necessary to limit the impact of the sparseness, or relatively low population, of some datasets, such as reflectivity (see the supporting information for more information). The linear correlation parameters between the threshold and other key parameters were calculated for each group. This includes material (from PL and lifetime measurements), cavity (dimensions and reflectivity) and performance (lasing wavelength) properties for each group, shown in Figure 3(a). This provides an at-a-glance means of assessing the performance, averaging over any inter-type variation in NW growths, and enables global trends to be observed that are related to the underlying physics and performance of NWLs. Whilst many of the correlation parameters reported are small (with ), this is, at least partially, due to the large degree of variation in multiple other parameters across the NW population. These correlations are statistically significant at the level and can therefore be described with confidence.
These results elucidate the importance of the laser cavity on all classes of NWLs. The NW length has a statistically significant negative correlation with threshold, which, as previously discussed, is due to a increase in the volume of gain material. This dictates a clear route towards optimising the threshold of all types of NWLs: minimise the length variation between each NW. A similar effect is also observed for increasing the diameter of “bulk” and excitonic NWLs, which is compounded by an additional increase in the optical confinement factor. However, no diameter correlation is observed for heterostructure samples: this suggests that the optical gain is decoupled from the NW diameter and confinement factor, instead depending on the QW-thickness and QW-lasing mode overlap 5.
The strength of the length correlation reduces from excitonic (r = -0.17), to “bulk” (r = -0.11) and to heterostructure (r = -0.05) lasers. This suggests that the length has different degree of impact on the threshold for different NW classes. A possible explanation for this is due to differences in distributed losses between the NW classes, which would partially counter-balance the increased gain. For example, for heterostructure NWLs, the overgrowth of the shells can lead to distortions in the NW cross-section shape and consequently enhance the optical losses 73.
The cavity reflectivity also plays a crucial role in the performance, which shows a negative correlation with threshold for “bulk” and heterostructure lasers, indicating that a higher reflectivity reduces the optical losses and, thus, the threshold. No trend is observed for the excitonic NWs, although this may be due to a small sample size (90 wires in total). Another route to reduce variation in NWL performance is therefore to optimise the end facets, either by optimising the NW transfer 65 or standardising the NW diameter during growth 50.
The material properties of the NWLs do not show consistent correlations with threshold. The PL FWHM, an often-used measure of the inhomogeneity in the material 6, has no statistically significant impact on the threshold. The carrier lifetime also shows no correlation for “bulk” NWLs, which suggests that, despite the variation in lifetime shown in Fig. 2(d), there is little significant change in the quantum efficiency of these classes across their populations. In contrast, the heterostructure and excitonic lasers show a negative correlation between lifetime and threshold. This correlation has been discussed in detail previously 15, and is attributed to a variation in the non-radiative recombination rate (and quantum efficiency) which is related to a change in defect density within the gain medium. For these samples, there remains scope to improve the lasing performance by improving the IQE of the active layer by optimising the growth conditions to reduce the defect density.
The PL peak energy in “bulk” and heterostructure NWs show a positive correlation with threshold. The mechanism for the energy variation is related to composition, strain or heterostructure fluctuations, depending on the sample, and this also leads to shifts in the laser gain spectrum. These correlations are a consequence of the strong negative correlations between lasing wavelength and threshold which have been previously attributed to reabsorption 1 and band-filling effects 15. Interestingly, the excitonic devices show an opposite negative correlation between PL peak energy and threshold. This is consistent with previous studies that demonstrate the importance of strong exciton binding to achieve lasing 74: NWs with more strongly bound excitons are more likely to have a lower lasing threshold. This is despite the lasing action being sustained by an electron-hole plasma which forms at high fluences 75.
The wavelength correlation for each NW class shows the same effects to a different degree. The excitonic devices, shown in Figure 3(b), have minimal reabsorption since the peak of the gain spectrum lies in the material bandgap, the 3D density of states and excitonic recombination also reduces the impact of band-filling and so the correlation between wavelength and threshold is the lowest of the three classes. The “bulk” class (Fig. 3(c)) has more reabsorption, and so the correlation is increased. Finally, the heterostructure NWs (Fig. 3(d)) exhibit reabsorption into core and cap layers, and will exhibit more band-filling due to a reduced 2D density of states 15 and thus have the highest observed correlation.
In summary, we have demonstrated a data-driven experimental approach to study large numbers of micron-scale opto-electronic devices. We have applied this to investigate the properties of 55,516 individual NWLs with nine different designs and have extracted parameters that describe how the gain medium, lasing cavity and performance vary across the population. We have shown how the approach can be used to effectively screen for the best-in-class devices and we have quantified the statistical distribution of each parameter and used this to compare the properties of different NWL types in a statistically robust manner. Through a correlative analysis, it was determined that the properties of the laser cavity are consistently a driving factor in the performance of NWLs and this elucidates a route towards achieving homogeneous performance across a population of any type of NWLs: by optimising the NW length and end-facet reflectivity.
Supporting Information
Supporting information is available for this paper, which includes the growth details for each type of nanowire, some further experimental details, low power PL spectra and statistics of the NW diameters and lifetimes. This can be found at DOI: ????.
Research data supporting this publication is available at DOI: 10.48420/26403496, and the code to perform the analysis will be made available at DOI: (DOI to be provide in due course), and on github: (URL to be provide in due course).
Acknowledgements
This work was funded by UKRI under grants MR/T021519/1, EP/V036343/1 and EP/W002302/1, by the Deutsche Forschungsgemeinschaft within the frame of the collaborative research center CRC 1375 (project C5), by the Australian Research Council, the Australian National Fabrication Facility ACT node, as well as China special Grant (134000-E62201ZJ); Zhejiang Provincial Natural Science Foundation of China (Z24F040009), Zhejiang University Education Foundation Qizhen Scholar Foundation (K20240015), and CMOS Special Program of Zhejiang University (04010000-K2A033208).
CRediT author statement: Stephen Church: Data curation, Formal analysis, Investigation, Methodology, Software, Visualisation, Writing - original draft. Francesco Vitale, Aswani Gopakumar, Nikita Gagrani,Yunyan Zhang, Nian Jiang, Hark Hoe Tan, Chennupati Jagadish, Huiyun Liu, Hannah Joyce and Carsten Ronning: Resources, Writing - review and editing. Patrick Parkinson: Conceptualization, Data curation, Funding acquisition, Methodology, Software, Supervision, Writing - review and editing.